OpenStudy (anonymous):

how do you solve for x? III. the quantity of 2x plus 1, divided by 3 equals the quantity of x plus 6, divided by 2

OpenStudy (marigirl):

cross multiply is a simple n straightforward one to do 

OpenStudy (marigirl):

4x+2=3x+18 4x-3x=18-2 x=16

OpenStudy (anonymous):

THANK YOU SOOOO MUCH

OpenStudy (marigirl):

sry couldnt help myself and went straight to the ans :D

OpenStudy (anonymous):

No, no that's fantastic lol I hate math and appreciate straight to the point answers :)